Round your answer to this question to the nearest year. Find the number of years it would take for R748,00

to increase by 72% if it is invested at an interest rate of 7,9% per annum, compounded semi-annually.


1
Expert's answer
2021-05-28T10:22:20-0400

"1.72\cdot748=748(1+0.079/2)^{2n}"

"1.72=(1+0.079\/2)^{2n}"

"n=ln1.72/(2ln(1+0.079/2))=7\ years"